# TIM: An Efficient Temporal Interaction Module for Spiking Transformer

### v1

@@ Abstract
Spiking Neural Networks (SNNs), as the third generation of neural networks, have gained prominence for their biological plausibility and computational efficiency, especially in processing diverse datasets. The integration of attention mechanisms, inspired by advancements in neural network architectures, has led to the development of Spiking Transformers. These have shown promise in enhancing SNNs’ capabilities, particularly in the realms of both static and neuromorphic datasets. Despite their progress, a discernible gap exists in these systems, specifically in the Spiking Self Attention (SSA) mechanism’s effectiveness in leveraging the temporal processing potential of SNNs. To address this, we introduce the Temporal Interaction Module (TIM), a novel, convolution-based enhancement designed to augment the temporal data processing abilities within SNN architectures. TIM’s integration into existing SNN frameworks is seamless and efficient, requiring minimal additional parameters while significantly boosting their temporal information handling capabilities. Through rigorous experimentation, TIM has demonstrated its effectiveness in exploiting temporal information, leading to state-of-the-art performance across various neuromorphic datasets.
@@ 1 Introduction
Spiking neural networks (SNNs), representing a novel paradigm in the evolution of artificial neural networks (ANNs), derive their foundational principles from the intricacies of biological neural systems, with a particular focus on dynamic and temporal processing capabilities Maass (1997); Zeng et al. (2023). The event-driven mechanism inherent in SNNs markedly amplifies their energy efficiency, while simultaneously fostering improved interoperability with a broad range of neuromorphic hardware Roy et al. (2019). These networks have emerged as a pivotal tool within the domain of cognitive and intelligent systems research.
@@
  • To address this, we have designed a Temporal Interaction Module. This module adaptively utilizes information from different time steps and functions as a lightweight addition. It can be integrated with existing attention computation modules without significantly increasing computational load.
  • We conducted experiments on neuromorphic datasets, including DVS-CIFAR10, NCALTECH101, NCARS, UCF101-DVS, and HMDB-DVS. Our results demonstrate that our approach sets a new benchmark in the Spiking Transformer domain, achieving optimal performance across these datasets.
@@ 2.1 Advances in SNN Temporal Processing
In terms of enhancing the temporal information processing ability of SNNs, significant progress has been made by researchers. For instance,  Kim et al. (2023) explored the dynamic characteristics of temporal information in SNNs by estimating the Fisher information of weights. PLIF Fang et al. (2021b) enhanced the integration of information at different time steps in SNNs by introducing a learnable membrane potential constant. IIRSNN Fang et al. (2020) improved the spatiotemporal information processing capabilities of SNNs through synaptic modeling. TA-SNN Yao et al. (2021) introduced a temporal-based attention mechanism, enabling the adaptive allocation of importance to different time steps.  Shen et al. (2024) significantly bolstered the temporal information processing capability of SNNs by incorporating dendritic nonlinear computations. ETC Zhao et al. (2023) introduced Temporal Enhancement Consistency constraints to enable SNNs to learn from outputs at different time steps, thereby enhancing performance and reducing latency. BioEvo Shen et al. (2023) employed a biologically inspired neural circuit search approach, adaptively coordinating different circuits to improve the performance of SNNs in perceptual tasks and reinforcement learning.
@@ 2.2 Spiking Transformer
In the Transformer domain, Spikformer Zhou et al. (2022) innovatively implemented an attention mechanism called Spiking Sequence Attention. This mechanism replaced the traditional activation function of ANNs with spiking neurons and omitted the softmax function typically required before calculating attention Vaswani et al. (2017), opting instead for direct matrix multiplication. DISTA Xu et al. (2023) enhanced the capture of spatiotemporal data by designing new neuronal connections. Spikeformer Li et al. (2022a) introduced separate Temporal and Spatial Multi-head Self Attention in each Transformer block to strengthen data processing capabilities. Spikingformer Zhou et al. (2023a) achieved greater efficiency by rearranging the positions of neurons and convolutional layers to eliminate float-integer multiplications. In contrast, the Spike-driven Transformer Yao et al. (2023a) reduced time complexity by altering the order of attention computation and substituting multiplication with addition.
@@ 3.1 Spiking Neuron
In our study, the variable $a$ serves as a hyperparameter, instrumental in dictating the configuration of the surrogate gradient’s shape. We have strategically set the value of $a$ to 2.
@@ 4.2 TIM Unit
Figure 2: Overview of the Temporal Interaction Module: Adaptive Integration of Historical and Current Temporal Data.
As shown in Fig. 2, our design incorporates a temporal interaction module within the query structure, aiming to better account for historical information, thereby enhancing the processing of temporal features. The details are elucidated in Eq. 7:
[…]
Here, $f$ signifies an operation for extracting historical information, implemented as a one-dimensional convolution in this study. Consequently, $Q^{TIM}[t]$ comprises two components: the first represents the contribution of historical information to the current attention, while the second signifies the direct impact of the current input. The hyperparameter $alpha$ facilitates an adaptive balance between the significance of historical and current information.
@@ 4.3 TIM Stream
As depicted in Fig. 3, this module is integrated into the computational graph of the attention matrix.
$A^{TIM}[t]$ $=A^{TIM}[t]K[t]^{T}V[t]$ (8)
$=\alpha f(Q^{TIM}[t-1])K[t]^{T}V[t]+(1-\alpha)Q[t]^{T}V[t]$
@@
Figure 3: The upper part of the figure describes the process of SSA, while the lower part illustrates the process after integrating TIM Stream into SSA.
@@ 5 Experiments
2D-WT-Haar Wang et al. (2023) SPikformer 16 81.0 - -
TIM Spikformer 10 81.6 79.0 96.5
Table 1: Comparison with other benchmark results on CIFAR10-DVS, N-CALTECH101 and NCARS
@@ 5.1 CIFAR10-DVS
DVS-CIFAR10 Li et al. (2017) is an event stream dataset comprising 10,000 images from the CIFAR-10 dataset. These images are converted into event streams using a bicubic interpolation method. These images are adeptly transformed into event streams through a bicubic interpolation method, where each image is represented as a sequence of events, each signifying a spike in pixel activation. As outlined in Tab. 1, our algorithm exhibits a significant improvement over the prevalent Spiking Transformer framework. When benchmarked against Spikformer, our Temporal Interaction Module demonstrates enhanced performance, exceeding it by approximately 2.7% under identical configurations. Notably, even when Spikformer is optimized with a longer time step (TimeStep=16), our algorithm maintains a performance lead of 0.7%. While Spikeformer attains an accuracy of 81.4% in 4 steps, it requires a substantial 9.28 million parameters. To our knowledge, this achievement marks the most advanced state-of-the-art (SOTA) result for Spiking Transformer applications on the CIFAR10-DVS dataset. In comparison with convolutional structures, the efficacy of TIM is on par with EventMix Shen et al. (2022b), which records a similar accuracy of 81.45%. However, a striking advantage of our model is its efficiency: while EventMix operates with an extensive 11.69 million parameters, our TIM model achieves comparable results with a significantly reduced parameter count of only 2.59 million.
@@ 5.2 N-CALTECH101
The N-CALTECH101 dataset, as introduced in Orchard et al. (2015), is an innovative adaptation of the classic CALTECH101 image library, covering 101 diverse categories. This dataset is notably derived using advanced neuromorphic visual sensors. It uniquely offers a time-based visual event representation of the original static image categories, effectively capturing the dynamic and temporal nuances inherent in each category. In our research, we have compiled and analyzed several benchmarks from the N-CALTECH101 dataset alongside the results of our Temporal Integration Model, as summarized in Tab. 1. These benchmarks primarily stem from ANN architectures. However, our SNN model based TIM exhibits a remarkable performance edge over these benchmarks. Notably, even when compared to the Event Transformer, which falls within the same Transformer category, our SNN-based TIM achieves a significant lead, outperforming it by 0.1%.
@@ 5.3 NCARS
The NCARS dataset, as introduced in Orchard et al. (2015), represents a cutting-edge binary classification dataset. It consists of a comprehensive collection of dynamic event streams, meticulously capturing cars and non-car objects through the lens of event cameras. These event streams offer a rich, real-time depiction of visual phenomena, distinguishing themselves by their ability to encode temporal information about moving objects. Our TIM continues to set new benchmarks in the analysis of the NCARS dataset. It achieves a remarkable 96.5% accuracy, a testament to its robustness and advanced feature extraction capabilities. This level of performance not only surpasses the Event Transformer by a notable margin of 1.1% but also maintains a significant lead over EventMix, outperforming it by 0.2%. These results underscore the efficacy of TIM in handling the complex dynamics and temporal variations inherent in the NCARS dataset.
@@
Table 2: Comparison with other benchmark results on UCF101-DVS and HMDB51-DVS
Figure 4: Temporal enhancement validation of TIM on CIFAR10-DVS and N-CALTECH101 and confusion matrix of TIM and Local TIM on CIFAR10-DVS.
@@ 5.4 UCF101-DVS And HMDB51-DVS
The UCF101-DVS and HMDB51-DVS datasets represent neuromorphic adaptations of the well-known UCF101 and HMDB51 datasets, respectively. These adaptations transform the original video-based datasets into a neuromorphic format, creating a rich collection of visual event streams. The UCF101-DVS encompasses an array of 101 action categories, while the HMDB51-DVS includes 51 distinct action categories. In these datasets, each action category is meticulously converted into dynamic visual event streams using advanced event cameras, offering a novel perspective on action recognition. Our TIM has showcased exceptional performance on these challenging datasets. On the HMDB51-DVS dataset, TIM achieved a top-1 accuracy of 58.6%, and on the UCF101-DVS dataset, it reached an even more impressive top-1 accuracy of 63.8%, as detailed in Tab. 2.
@@
Baseline 78.5
CIFAR10-DVS Baseline + TIM 81.6
Baseline 77.7
N-CHALTECH101 Baseline + TIM 79.0
@@ 6.2 Temporal Enhancement Validation
To rigorously validate that the performance enhancements observed in the Spiking Transformer are indeed due to its improved temporal processing capabilities, and not merely a consequence of an increased parameter count, we undertook a meticulous adjustment of the Temporal Interaction Module’s structure. This refinement was crucial in isolating the impact of temporal capability enhancements from the effects of parameter augmentation. As detailed in Eq. (9) and (10), we specifically re-engineered the operational mechanism of the TIM Stream. This modification was aimed at optimizing the functional dynamics of the TIM Unit. Instead of allowing the TIM Unit to engage in interactions across multiple time steps, which could potentially confound our assessment of its temporal processing proficiency, we constrained its operation to the current time step only (which we call local TIM).
@@
$attention=Q[t]K^{T}[t]V[t]$ (10)
As illustrated in Fig. 5, a notable experiment was conducted applying our model to the CIFAR10-DVS and N-CALTECH101 datasets. The results were quite revealing: the model’s accuracy decreased by 2.7% and 2.4% on these datasets, respectively. This phenomenon strongly indicates that the improvement in accuracy previously attributed to the Temporal Interaction Module is primarily a result of its enhanced ability to process temporal information, rather than being a consequence of an increased parameter count. The stability of the parameter count during the experiment reinforces this conclusion, highlighting the intrinsic value of TIM in specifically enhancing temporal data processing capabilities. The comprehensive training details, along with the corresponding confusion matrix, are presented in Fig. 4.
Figure 5: Temporal Enhancement Validation on CIFAR10-DVS and N-CALTECH101 datasets.
We further concentrated on investigating the impact of different values of the hyperparameter $\alpha$ on the performance of the Temporal Integration Module. This exploration was conducted through a series of experiments using the CIFAR10-DVS and N-CALTECH101 datasets. The findings from these experiments were quite revealing. We observed that regardless of the specific value, setting $\alpha$ to any non-zero number consistently resulted in better performance compared to the baseline scenario where $\alpha$ was set to zero. This pattern indicates that the introduction of temporal interaction in TIM significantly enhances its performance. More specifically, for the CIFAR10-DVS dataset, the model’s peak performance, with an accuracy of 81.6%, was achieved when $\alpha$ was set to 0.6. Similarly, for the N-CALTECH101 dataset, the optimal performance occurred at an $\alpha$ value of 0.4, reaching a top accuracy of 79.0%. These results highlight an important aspect of TIM’s functionality. While the model exhibits robustness to a range of $\alpha$ values, fine-tuning this hyperparameter allows for more precise optimization relative to specific datasets. The ability to adjust $\alpha$ effectively tailors the model to different data distributions, enhancing TIM’s versatility and efficacy in diverse neuromorphic data processing applications.
Figure 6: The top-1 accuracy of N-CALTECH101 and CIFAR10-DVS with different alpha values.

### v3

@@ Abstract
Spiking Neural Networks (SNNs), as the third generation of neural networks, have gained prominence for their biological plausibility and computational efficiency, especially in processing diverse datasets. The integration of attention mechanisms, inspired by advancements in neural network architectures, has led to the development of Spiking Transformers. These have shown promise in enhancing SNNs’ capabilities, particularly in the realms of both static and neuromorphic datasets. Despite their progress, a discernible gap exists in these systems, specifically in the Spiking Self Attention (SSA) mechanism’s effectiveness in leveraging the temporal processing potential of SNNs. To address this, we introduce the Temporal Interaction Module (TIM), a novel, convolution-based enhancement designed to augment the temporal data processing abilities within SNN architectures. TIM’s integration into existing SNN frameworks is seamless and efficient, requiring minimal additional parameters while significantly boosting their temporal information handling capabilities. Through rigorous experimentation, TIM has demonstrated its effectiveness in exploiting temporal information, leading to state-of-the-art performance across various neuromorphic datasets. The code is available at https://github.com/BrainCog-X/Brain-Cog/tree/main/examples/TIM.
@@ 1 Introduction
Spiking neural networks (SNNs), representing a novel paradigm in the evolution of artificial neural networks (ANNs), derive their foundational principles from the intricacies of biological neural systems, with a particular focus on dynamic and temporal processing capabilities Maass (1997); Zeng et al. (2023). The event-driven mechanism inherent in SNNs markedly amplifies their energy efficiency, while simultaneously fostering improved interoperability with a broad range of neuromorphic hardware Roy et al. (2019); Wei et al. (2024). These networks have emerged as a pivotal tool within the domain of cognitive and intelligent systems research.
@@
  • To address this, we have designed a Temporal Interaction Module to adaptively utilizes information from different time steps and functions as a lightweight addition. It can be integrated with existing attention computation modules without significantly increasing computational load.
  • We conducted experiments on neuromorphic datasets, including CIFAR10-DVS, NCALTECH101, NCARS, UCF101-DVS, HMDB-DVS and SHD. We demonstrated the effectiveness and generalization ability of our method. Our results show that our approach sets a new benchmark in the Spiking Transformer domain, achieving better performance across these datasets.
@@ 2.1 Advances in SNN Temporal Processing
In terms of enhancing the temporal information processing ability of SNNs, significant progress has been made by researchers. For instance,  Kim et al. (2023) explored the dynamic characteristics of temporal information in SNNs by estimating the Fisher information of weights. PLIF Fang et al. (2021b) enhanced the integration of information at different time steps in SNNs by introducing a learnable membrane potential constant. IIRSNN Fang et al. (2020) improved the spatiotemporal information processing capabilities of SNNs through synaptic modeling. Zhang et al. (2021) discovered that learning based on spike timing in an event-driven manner can yield significant improvements in classification accuracy.TA-SNN Yao et al. (2021) introduced a temporal-based attention mechanism, enabling the adaptive allocation of importance to different time steps. TCJA Zhu et al. (2024) incorporates an attention mechanism designed to assess the significance of pulses in both temporal and spatial dimensions. Shen et al. (2024) significantly bolstered the temporal information processing capability of SNNs by incorporating dendritic nonlinear computations. ETC Zhao et al. (2023) introduced Temporal Enhancement Consistency constraints to enable SNNs to learn from outputs at different time steps, thereby enhancing performance and reducing latency. BioEvo Shen et al. (2023) employed a biologically inspired neural circuit search approach, adaptively coordinating different circuits to improve the performance of SNNs in perceptual tasks and reinforcement learning.
@@ 2.2 Spiking Transformer
In the Transformer domain, Spikformer Zhou et al. (2022) innovatively implemented an attention mechanism called Spiking Self Attention. This mechanism replaced the traditional activation function of ANNs with spiking neurons and omitted the softmax function typically required before calculating attention Vaswani et al. (2017), opting instead for direct matrix multiplication. DISTA Xu et al. (2023) enhanced the capture of spatiotemporal data by designing new neuronal connections. Spikeformer Li et al. (2022a) introduced separate Temporal and Spatial Multi-head Self Attention in each Transformer block to strengthen data processing capabilities. Spikingformer Zhou et al. (2023a) achieved greater efficiency by rearranging the positions of neurons and convolutional layers to eliminate float-integer multiplications. In contrast, the Spike-driven Transformer Yao et al. (2023a) reduced time complexity by altering the order of attention computation and substituting multiplication with addition.
@@ 3.1 Spiking Neuron
In our study, the variable $a$ serves as a hyperparameter, instrumental in dictating the configuration of the surrogate gradient’s shape. We have strategically set the value of $a$ to 4.
@@ 3.2 Spiking Self Attention Analysis
Figure 2: (a): Overview of the Temporal Interaction Module: Adaptive Integration of Historical and Current Temporal Data. (b): The upper part of the figure describes the process of SSA, while the lower part illustrates the process after integrating TIM Stream into SSA.
@@ 4.2 TIM Unit
Here, $f$ signifies an operation for extracting historical information, implemented as a one-dimensional convolution in this study. Consequently, $Q^{TIM}[t]$ comprises two components: the first represents the contribution of historical information to the current attention, while the second signifies the direct impact of the current input. The hyperparameter $\alpha$ facilitates an adaptive balance between the significance of historical and current information. The specific illustration of TIM Unit can be found in Fig 2(a).
@@ 4.3 TIM Stream
As depicted in Fig. 2(b), this module is integrated into the computational graph of the attention matrix.
$Attention[t]=A^{TIM}[t]K[t]^{T}V[t]$ (7)
$=\alpha f(Q^{TIM}[t-1])K[t]^{T}V[t]+(1-\alpha)Q[t]K[t]^{T}V[t]$
@@ 5 Experiments
TIM(Ours) Spikformer 10 81.6 79.0 96.5
Table 1: Comparison with other benchmark results on CIFAR10-DVS, N-CALTECH101 and NCARS. The * indicates unknown time steps from the original studies.
@@ 5.1 CIFAR10-DVS
CIFAR10-DVS Li et al. (2017) is an event stream dataset comprising 10,000 images from the CIFAR-10 dataset. These images are converted into event streams using a bicubic interpolation method. As outlined in Tab. 1, our algorithm exhibits a significant improvement over the prevalent Spiking Transformer framework. When benchmarked against Spikformer, our TIM demonstrates enhanced performance, exceeding it by approximately 2.7% under identical configurations. Notably, even when Spikformer is optimized with a longer time step (Time Step=16), our algorithm maintains a performance lead of 0.7%. While Spikeformer attains an accuracy of 81.4% in 4 steps, it requires a substantial 9.28 million parameters. To our knowledge, this achievement marks the most advanced state-of-the-art (SOTA) result for Spiking Transformer applications on the CIFAR10-DVS dataset. In comparison with convolutional structures, the efficacy of TIM is on par with EventMix Shen et al. (2022b), which records a similar accuracy of 81.45%. However, a striking advantage of our model is its efficiency: while EventMix operates with an extensive 11.69 million parameters, our TIM model achieves comparable results with a significantly reduced parameter count of only 2.59 million.
@@ 5.2 N-CALTECH101
The N-CALTECH101 dataset, as introduced in Orchard et al. (2015), is an innovative adaptation of the classic CALTECH101 image library, covering 101 diverse categories. This dataset is notably derived using advanced neuromorphic visual sensors. It uniquely offers a time-based visual event representation of the original static image categories, effectively capturing the dynamic and temporal nuances inherent in each category. In our research, we have compiled and analyzed several benchmarks from the N-CALTECH101 dataset alongside the results of our Temporal Integration Model, as summarized in Tab. 1. These benchmarks primarily stem from ANN architectures. However, our SNN model based TIM exhibits a remarkable performance edge over these benchmarks. Despite being compared with the Event Transformer which is an ANN-based Transformer model, our SNN-based TIM has achieved comparable performance.
@@ 5.3 NCARS
The NCARS dataset, as introduced in Orchard et al. (2015), represents a cutting-edge binary classification dataset. It consists of a comprehensive collection of dynamic event streams, meticulously capturing cars and non-car objects through the lens of event cameras. These event streams offer a rich, real-time depiction of visual phenomena, distinguishing themselves by their ability to encode temporal information about moving objects. Our TIM continues to set new benchmarks in the analysis of the NCARS dataset. It achieves a remarkable 96.5% accuracy, a testament to its robustness and advanced feature extraction capabilities. This level of performance not only surpasses the Event Transformer by a notable margin of 1.1% but also maintains a weak lead over EventMix, outperforming it by 0.2%. These results underscore the efficacy of TIM in handling the complex dynamics and temporal variations inherent in the NCARS dataset.
@@
SHD Spikformer 85.1
TIM(Ours) 86.3
Table 2: Comparison with other benchmark results on UCF101-DVS, HMDB51-DVS and SHD.
@@ 5.4 UCF101-DVS And HMDB51-DVS
The UCF101-DVS and HMDB51-DVS datasets represent neuromorphic adaptations of the well-known UCF101 and HMDB51 datasets, respectively. These adaptations transform the original video-based datasets into a neuromorphic format, creating a rich collection of visual event streams. The UCF101-DVS encompasses an array of 101 action categories, while the HMDB51-DVS includes 51 distinct action categories. In these datasets, each action category is meticulously converted into dynamic visual event streams using advanced event cameras, offering a novel perspective on action recognition. Our TIM has showcased exceptional performance on these challenging datasets. On the HMDB51-DVS dataset, TIM achieved a top-1 accuracy of 58.6%, and on the UCF101-DVS dataset, it reached a top-1 accuracy of 63.8%, as detailed in Tab. 2.
@@
CIFAR10-DVS Baseline 78.5
Baseline + TIM 81.6
N-CALTECH101 Baseline 77.7
Baseline + TIM 79.0
@@
5.5 SHD
The Spiking Heidelberg Digits (SHD) dataset is an audio-based classification dataset containing 1000 spoken digits ranging from zero to nine in both English and German languages, with a total of 20 classes. Speech datasets typically exhibit more complex interdependence between time steps. TIM demonstrates a 1.2% advantage over the baseline model on this dataset, shown in Tab 2. The results indicates TIM’s capability to extract intricate temporal features.
@@ 6 Discussion
Figure 3: Temporal enhancement validation of TIM on CIFAR10-DVS and N-CALTECH101.
@@ 6.2 Temporal Enhancement Validation
To rigorously validate that the performance enhancements observed in the Spiking Transformer are indeed due to its improved temporal processing capabilities, and not merely a consequence of an increased parameter count, we undertook a meticulous adjustment of the Temporal Interaction Module’s structure. This refinement was crucial in isolating the impact of temporal capability enhancements from the effects of parameter augmentation. As detailed in Eq. (8) and (9), we specifically re-engineered the operational mechanism of the TIM Stream. This modification was aimed at changing the functional dynamics of the TIM Unit. Instead of allowing the TIM Unit to engage in interactions across multiple time steps, which could potentially confound our assessment of its temporal processing proficiency, we constrained its operation to the current time step only (which we call local TIM).
@@
$Attention[t]=Q[t]K^{T}[t]V[t]$ (9)
As illustrated in Tab. 4, a notable experiment was conducted applying our model to the CIFAR10-DVS and N-CALTECH101 datasets. The results were quite revealing: the model’s accuracy decreased by 2.7% and 2.4% on these datasets, respectively. This phenomenon strongly indicates that the improvement in accuracy previously attributed to the Temporal Interaction Module is primarily a result of its enhanced ability to process temporal information, rather than the increased parameter count. The stability of the parameter count during the experiment reinforces this conclusion, highlighting the intrinsic value of TIM in specifically enhancing temporal data processing capabilities. The comprehensive training details are presented in Fig. 3.
Dataset Model Acc@1(%)
CIFAR10-DVS Baseline+local TIM 78.9
Baseline + TIM 81.6
N-CALTECH101 Baseline+local TIM 76.6
Baseline + TIM 79.0
Table 4: Temporal Enhancement Validation on CIFAR10-DVS and N-CALTECH101 datasets.
alpha 0 0.2 0.4 0.5 0.6 0.8
CIFAR10-DVS 78.5 79.8 79.7 80.7 81.6 81.2
N-CALTECH101 77.7 78.5 79.0 78.5 77.6 78.7
Table 5: The top-1 accuracy of N-CALTECH101 and CIFAR10-DVS with different alpha values.
We further concentrated on investigating the impact of the hyperparameter $\alpha$ on the performance of the Temporal Integration Module. This exploration was conducted through a series of experiments using the CIFAR10-DVS and N-CALTECH101 datasets. The top-1 accuracy of TIM on both datasets with different $\alpha$ values are illustrated in Tab. 5. The findings from these experiments were quite revealing. We observed that regardless of the specific value, setting $\alpha$ to any non-zero number consistently resulted in better performance compared to the baseline scenario where $\alpha$ was set to zero. This pattern indicates that the introduction of temporal interaction in TIM significantly enhances its performance. More specifically, for the CIFAR10-DVS dataset, the model’s peak performance, with an accuracy of 81.6%, was achieved when $\alpha$ was set to 0.6. Similarly, for the N-CALTECH101 dataset, the optimal performance occurred at an $\alpha$ value of 0.4, reaching a top accuracy of 79.0%. These results highlight an important aspect of TIM’s functionality. While the model exhibits robustness to a range of $\alpha$ values, fine-tuning this hyperparameter allows for more precise optimization relative to specific datasets. The ability to adjust $\alpha$ effectively tailors the model to different data distributions, enhancing TIM’s versatility and efficacy in diverse neuromorphic data processing applications.
6.3 Efficiency and Generalizability Validation
Figure 4: Efficiency and Generalizability Validation of TIM on CIFAR10-DVS.
TIM introduces additional operations into the model, resulting in extra computational overhead. To alleviate concerns about TIM’s reduced efficiency, we reduced the training steps. We found that with 6 steps, TIM achieves the same performance as the baseline does with 10 steps, shown in Fig 4. This implies that TIM can save approximately 40% of the time, suggesting that TIM retains the efficiency of SNN.
[…]
SDSA achieved a top-1 accuracy of 77% on CIFAR10-DVS with 10 steps, while SDSA+TIM achieved a top-1 accuracy of 78.5%. We consider the 1.5% difference to be statistically significant, indicating that TIM gains traction in SDSA, demonstrating its ability to generalize across a broader range of Spiking Transformer architectures.